OpenStudy (anonymous):

Sn=Suumation from i=2 to inf = 3/4 -1/2n -1/2(n+1) who it happen?

OpenStudy (anonymous):

i am a bit lost on this one

OpenStudy (anonymous):

sorry typing mistake * how it happen?

OpenStudy (anonymous):

because i can't really read it but it looks like you have a telescoping sum write out the first few terms and see if that is true

OpenStudy (anonymous):

how we found the formula of this partial sum?

OpenStudy (anonymous):

can you try in equation editor? i can only make a guess as to what this says

OpenStudy (anonymous):

looks like maybe \[\frac{3}{4}-\frac{1}{2n}-\frac{1}{2(n+1)}\]

OpenStudy (anonymous):

\[Sn = \sum_{i=2}^{\infty} 1/i^2 -1 = 3/4 - 1/2n -1/2(n+1)\]

OpenStudy (anonymous):

1/(i^2-1)

OpenStudy (anonymous):

satelite you are right

OpenStudy (anonymous):

oh ok you are looking for \[\sum_i \frac{1}{i^2-1}\]

OpenStudy (anonymous):

how did we find this formula

OpenStudy (anonymous):

i bet we can use partial fractions for this one

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

how did 3/4 -1/2n -1/2(n+1) come?

OpenStudy (anonymous):

is it a general formula for this series?

OpenStudy (anonymous):

typo there \[\frac{1}{i^2-1}=\frac{1}{2(i-1)}-\frac{1}{2(i+1)}\]

OpenStudy (anonymous):

ok then?

OpenStudy (anonymous):

now imagine what happens if you start replacing i by 2, 3, 4, ...

OpenStudy (anonymous):

don't simplify, just replace

OpenStudy (anonymous):

how can we break it like this 1/(2(i-1) -1/2(i+1) will give us the denominator 2(i^2-1) whereas the original denominator is i^2-1

OpenStudy (anonymous):

ok that is a separate question lets take the first question first, that is, assume for a moment that we know \[\frac{1}{i^2-1}=\frac{1}{2(i-1)}-\frac{1}{2(i+1)}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and start replacing i by 2, 3, 4, 5, ... and see what happens you get \[\frac{1}{2}-\frac{1}{6}+\frac{1}{4}-\frac{1}{8}+\frac{1}{6}-\frac{1}{10}+\frac{1}{8}-\frac{1}{12}...\]

OpenStudy (anonymous):

with a little practice you will see this right away. the sum "telescopes" that is, you are adding and then subtracting the same terms, so they will end up canceling

OpenStudy (anonymous):

i.e. add up to zero

OpenStudy (anonymous):

okkkkk gotcha thanks bro

OpenStudy (turingtest):

why does the sum go to infinity and have an answer in terms of n ? me thinks we have a few typos here...

OpenStudy (anonymous):

that is, where you have \[\frac{1}{6}\] there will be a \[-\frac{1}{6}\] and then \[\frac{1}{8}\] and a \[-\frac{1}{8}\] and so on so what is left? the only numbers that do not have a "match" that is, the only numbers that remain are \[\frac{1}{2}+\frac{1}{4}=\frac{3}{4}\] which explains why the first term is \[\frac{3}{4}\]

OpenStudy (anonymous):

now to be pricise, we need to write out what \[\sum_{i=2}^n\frac{1}{i^2-1}\] is in terms of n

OpenStudy (anonymous):

the limit is zero so the sum won't be infinite

OpenStudy (anonymous):

isn't? @ satelite

OpenStudy (anonymous):

and from the telescoping nature of the sum, we see that what you will have is \[\frac{3}{4}-\frac{1}{2n}-\frac{1}{n(n+1)}\]

OpenStudy (anonymous):

the last two terms didn't get canceled yet, because we only went up to n now take the limit as n goes to infinity and you get \[\sum_{i=1}^{\infty}\frac{1}{i^2-1}=\lim_{n\to \infty}\sum_{i=2}^n\frac{1}{i^2-1}=\frac{3}{4}\]

OpenStudy (anonymous):

now as to how you get \[\frac{i}{i^2-1}=\frac{1}{2(i-1)}-\frac{1}{2(i+1)}\] do you know "partial fraction decomposition"? because that is what you need for this one

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

but have you read what I said just above?

OpenStudy (anonymous):

then it should be good from here yes?

OpenStudy (anonymous):

this line the limit is zero so the sum won't be infinite

OpenStudy (anonymous):

the limit of the second two terms is zero, so the limit of the whole thing is 3/4

OpenStudy (anonymous):

yes exactlly

OpenStudy (anonymous):

because that is the precise definition of the infinite sum, namely \[\sum_{i=2}^{\infty} a_i=\lim_{n\to \infty}\sum_{i=2}^na_i\] so you have to find the partial sums, and then compute the limit of the partial sums i would like to add that most people would say "the sum telescopes" and the only thing that survives is 3/4 and leave it at that, but it is good to be precise

OpenStudy (anonymous):

ok got it now I have lil' seperate question but related to series

OpenStudy (anonymous):

do me a favor and repost it is hard to keep scrolling down

OpenStudy (anonymous):

ok